A chemist obtain 18 grams of 10% saline solution from a mixture of 5% saline solution and a 15% saline solution. How much of the 15% saline solution is used in the mixture?
Answer by checkley77(12844)

.15X+.05(18-X)=.10*18
.15X+.9-.05X=1.8
.10X=1.8-.9
.10X=.9
X=.9/.10
X=9 GRAMS OF 15% SOLTION IS USED.
PROOF:
.15*9+.05(18-9)=.10*18
1.35+.05*9=1.8
1.35+.45=1.8
1.8=1.8